The product of two consecutive even integers is 12 less than 6 times their sum.
Answer by mananth(16946)

let the integerrs be x & x+2
Product = x(x+2)
12 less than 6 times the sum
6(x+x+2)-12
6(2x+2)-12

x(x+2)=6(2x+2)-12
x^2+2x=12x+12-12
x^2+2x-12x=0
x^2-10x=0
x(x-10)=0
x=0 OR 10
10,12